OpenStudy (anonymous):

Applications of derivative/ Find two numbers whose product is -12 and the sum of whose squares is a minimum.

OpenStudy (anonymous):

限制

OpenStudy (anonymous):

xy=-12 ==> y=-12/x ===> g(x) =x^2+(144/x^2) ==> g'(x) =2x-288/x^3 ===> g'(x) =0